A simple random sample with n=56 provided a sample mean of 21.0 and a sample standard deviation of 4.7 .

a. Develop a 90% confidence interval for the population mean (to 1 decimal).

b. Develop a 95% confidence interval for the population mean (to 1 decimal).

c.Develop a 99% confidence interval for the population mean (to 1 decimal).

d. What happens to the margin of error and the confidence interval as the confidence level is increased?
Answer #1

Solution :

Given that,

Point estimate = sample mean = \bar x = 21

sample standard deviation = s = 4.7

sample size = n = 56

Degrees of freedom = df = n - 1 = 55

a)

At 90% confidence level the t is ,

t\alpha /2,df = t0.05,55 = 1.673

Margin of error = E = t\alpha/2,df * (s /\sqrtn)

= 1.673 * (4.7 / \sqrt 56)

= 1.05

The 90% confidence interval estimate of the population mean is,

\bar x - E < \mu < \bar x + E

21 - 1.05 < \mu < 21 + 1.05

20.0 < \mu < 22.1

( 20.0 , 22.1 )

b)

At 95% confidence level the t is ,

t\alpha /2,df = t0.025,55 = 2.004

Margin of error = E = t\alpha/2,df * (s /\sqrtn)

= 2.004 * (4.7 / \sqrt 56)

= 1.2

The 95% confidence interval estimate of the population mean is,

\bar x - E < \mu < \bar x + E

21 - 1.2 < \mu < 21 + 1.2

19.8 < \mu < 22.2

( 19.8 , 22.8 )

c)

At 99% confidence level the t is ,

t\alpha /2,df = t0.005,55 = 2.668

Margin of error = E = t\alpha/2,df * (s /\sqrtn)

= 2.668 * (4.7 / \sqrt 56)

= 1.7

The 99% confidence interval estimate of the population mean is,

\bar x - E < \mu < \bar x + E

21 - 1.7 < \mu < 21 + 1.7

19.3 < \mu < 22.7

( 19.3 , 22.7 )

d)

The level of confidence increases , then the margin of error is increased and the confidence interval is wider .